if f is a linear function and f(0)=10 and f(10)=14, find f(5) and f(-5).
Answer by stanbon(75887)

if f is a linear function and f(0)=10 and f(10)=14, find f(5) and f(-5).
------------
f(0) gives you the y-intercept.
----------
Find the slope:
slope = (14-10)/(10-0) = 4/10 = 2/5
----------------------
EQUATION:
f(x) = (2/5)x + 10
--------------
f(5) = (2/5)*5 + 10 = 12
f(-5) = (2/5)*(-5) + 10 = 8
==============================
Cheers,
Stan H.